OpenStudy (anonymous):

evaluate using l'hopital i know you can't use l'hopital unless it is in the form of 0/0 or inf/inf. how can i transform this into one of those forms

OpenStudy (anonymous):

\[\lim_{x \rightarrow (\pi/2)+} \tan x- \sec x\]

OpenStudy (lgbasallote):

\[\tan x - \sec x \implies \frac{\sin x}{\cos x} -\frac{1}{\cos x} \implies \frac{\sin x - 1}{\cos x}\] does that help?

OpenStudy (dumbcow):

that should work

OpenStudy (anonymous):

yes soo much. So whenever I have a function such as this, the best thing to do is to convert to sin and cos?

OpenStudy (lgbasallote):

yup. always

OpenStudy (anonymous):

thanks, appreciate it!

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

the final answer is 0, just checking?

OpenStudy (lgbasallote):

why do you say so?

OpenStudy (anonymous):

because after finding the derivatives of num I got 0 and den i got -1. 0/-1 = 0

OpenStudy (lgbasallote):

but before you derive you need to verify it's 0/0 or infinity/infinity right? is this in that form?